# Any2Caption [Uncaptioned image]: Interpreting Any Condition to Caption for Controllable Video Generation

### v1

@@ Abstract
$\ast$\$$\ast$\$footnotetext: Work done during internship at Kuaishou Technology.🖂🖂footnotetext: Corresponding Author.
@@ 1 Introduction
Video serves as a fundamental medium for capturing real-world dynamics, making diverse and controllable video generation a key capability for modern artificial intelligence (AI) systems. Recently, video generation has gained significant attention, driven by advancements in Diffusion Transformers (DiT) [76, 29, 52, 44, 2], which have demonstrated the ability to generate realistic, long-duration videos from text prompts. These advancements have even led to industrial applications, such as filmmaking. However, we observe that a major bottleneck in the current video generation community lies in accurately interpreting user intention, so as to produce high-quality, controllable videos.
@@
In text-to-video (T2V) generation, studies [67, 30, 24] have suggested that detailed prompts, specifying objects, actions, attributes, poses, camera movements, and style, significantly enhance both controllability and video quality. Thus, a series of works have explored video recaption techniques (e.g., ShareGPT4Video [10], MiraData [30], and InstanceCap [15]) to build dense structured captions for optimizing generative models. While dense captions are used during training, in real-world inference scenarios, users most likely provide concise or straightforward input prompts [15]. Such a gap inevitably weakens instruction following and leads to suboptimal generation due to an incomplete understanding of user intent. To combat this, there are two possible solutions, manual prompt refinement or automatic prompt enrichment [15, 67] using large language models (LLMs). Yet, these approaches either require substantial human effort or risk introducing noise from incorrect prompt interpretations. As a result, this limitation in precisely interpreting user intent hinders the adoption of controllable video generation for demanding applications such as anime creation and filmmaking.
@@
In addition, to achieve more fine-grained controllable video generation, one effective strategy is to provide additional visual conditions besides text prompts—such as reference images [62, 17], identity [69, 22, 46], style [68, 42], human pose [45, 33], or camera [21, 75]—or even combinations of multiple conditions together [74, 41, 58]. This multimodal conditioning approach aligns well with real-world scenarios, as users quite prefer interactive ways to articulate their creative intent. Several studies have examined video generation under various conditions, such as VideoComposer [58], Ctrl-Adapter [41], and ControlVideo [74]. Unfortunately, these methods tend to rely on the internal encoders of Diffusion/DiT to parse rich heterogeneous input conditions with intricate requirements (e.g., multiple object IDs, and complex camera movements). Before generation, the model must accurately interpret the semantics of varied visual conditions in tandem with textual prompts. Yet even state-of-the-art (SoTA) DiT backbones have limited capacity for reasoning across different input modalities, resulting in suboptimal generation quality.
@@
To facilitate the any-to-caption instruction tuning for Any2Caption, we construct Any2CapIns, a large-scale dataset that converts a concise user prompt and diverse non-text conditions into detailed, structured captions. Concretely, the dataset encompasses four main categories of conditions: depth maps, multiple identities, human poses, and camera poses. Through extensive manual labeling combined with automated annotation by GPT-4V [1], followed by rigorous human verification, we curate a total of 337K high-quality instances, with 407K condition annotations, with the short prompts and structured captions averaging 55 and 231 words, respectively. In addition, we devise a comprehensive evaluation strategy to thoroughly measure the model’s capacity for interpreting user intent under these various conditions.
@@ 2 Related Work
Controllable video generation [53, 11, 16, 21] has long been a central topic in AI. Recent advanced DiT methods, such as OpenAI’s Sora [3] and HunyuanVideo [34], yield photorealistic videos over extended durations. Early work focused on text-controlled video generation [52, 24], the prevalent approach. Yet, text prompts alone may insufficiently capture user intent, spurring exploration of additional inputs including static images [62, 17], sketches [74, 58], human poses [77, 45, 33], camera views [75, 21], and even extra videos [32, 14, 70]. Thus, unifying these diverse conditions into an “any-condition” framework is highly valuable.
Recent works such as VideoComposer [58], Ctrl-Adapter [41], and ControlVideo [74] have explored any-condition video generation. However, they face challenges in controlling multiple modalities due to the limited interpretability of text encoders in Diffusion or DiT. Motivated by existing MLLMs’ multimodal reasoning [43, 39, 56], we propose leveraging an MLLM to consolidate all possible conditions into structured dense captions for better controllable generation. SoTA DiT models already exhibit the capacity to interpret dense textual descriptions as long as the input captions are sufficiently detailed in depicting both the scene and the intended generation goals. Thus, our MLLM-based encoder alleviates the comprehension bottleneck, enabling higher-quality video generation. To our knowledge, this is the first attempt in the field of any-condition video generation. Moreover, as the captioning stage is decoupled from backbone DiT, Any2Caption can integrate with existing video generation solutions without additional retraining.
Our approach also relates to video recaptioning, as our system produces dense captions from given conditions. In text-to-video settings, prior work [15, 67, 47, 27] shows that recaptioning yields detailed annotations that improve DiT training. For instance, ShareGPT4Video [10] uses GPT-4V [1] to reinterpret video content, while MiraData [30] and InstanceCap [15] focus on structured and instance-consistent recaptioning. Unlike these methods, we avoid retraining powerful DiT models with dense captions by training an MLLM as an any-condition encoder on pairs of short, dense captions that are easier to obtain. Moreover, recaptioning entire videos can introduce noise or hallucinations that undermine DiT training, whereas our framework sidesteps this risk. Finally, while previous studies rely on dense-caption-trained DiT models, the real-world user concise prompts might create a mismatch that degrades generation quality.
@@ 3 Any2CapIns Dataset Construction
While relevant studies recaption target videos for dense captions for enhanced T2V generation [10, 15, 30], these datasets exhibit two key limitations: (1) the absence of non-text conditions, and (2) short prompts that do not account for interactions among non-text conditions, potentially leading to discrepancies in real-world applications. To address these limitations, we introduce a new dataset, Any2CapIns, specifically designed to incorporate diverse multimodal conditions for generating structured video captions. The dataset is constructed through a three-step process (cf. Fig. 2), including data collection, structured caption construction, and user-centric short prompt generation.
@@
We begin by systematically categorizing conditions into four primary types: 1) Spatial-wise conditions, which focus on the structural and spatial properties of the video, e.g., depth maps, sketches, and video frames. 2) Action-wise conditions, which emphasize motion and human dynamics in the target video, e.g., human pose, motion. 3) Composition-wise conditions, which focus on scene composition, particularly in terms of object interactions and multiple identities in the target video. 4) Camera-wise conditions, which control video generation from a cinematographic perspective, e.g., camera angles, movement trajectories. Since it is infeasible to encompass all possible conditions in dataset collection, we curate representative datasets under each category, specifically including depth maps, human pose, multiple identities, and camera motion. During the data collection process, we leverage SoTA tools to construct conditions. For instance, Depth Anything [65] is used to generate depth maps, DW-Pose [66] provides human pose annotations, and SAM2 [50] is utilized for segmentation construction. In total, we collect 337K video instances and 407K conditions, with detailed statistics of the dataset presented in Tab. 1.
@@
The granularity of a caption, specifically the detailed selection of elements they encompass, plays a critical role in guiding the model to produce videos that closely align with the users’ intentions while preserving coherence and realism. Drawing inspiration from [30], we inherit its structured caption format consisting of (1) Dense caption, (2) Main object caption, (3) Background caption, (4) Camera caption, and (5) Style caption. Furthermore, the action descriptions of the subjects significantly influence the motion smoothness of the videos [57], we explicitly incorporate the (6) Action caption of subjects to form the final structured caption. Technically, following [57], we independently generate each caption component and subsequently integrate them to construct the final structured caption.
@@ 4 Any2Caption Model
Similar to existing MLLMs [43, 63, 39], Any2Caption incorporates an image encoder $F_{\mathcal{I}}$, a video encoder $F_{\mathcal{V}}$, a motion encoder $F_{\mathcal{M}}$ and a camera encoder $F_{\mathcal{C}}\}$ to process non-text conditions. These encoders are then integrated into an LLM backbone $F_{\mathcal{LLM}}$ (i.e., Qwen2-LLM) to facilitate structured video captioning. Specifically, we leverage a ViT-based visual encoder from Qwen2-VL as $F_{\mathcal{I}}$ and $F_{\mathcal{V}}$ for the unified modeling of images and videos, achieving effective interpretation of input conditions represented in image or video formats, such as depth maps and multiple identities. To enable human pose understanding, we represent the extracted human pose trajectories as $\bm{H}$=$\{(x_{n}^{k},y_{n}^{k})|k=1,\cdots,K,n=1,\cdots,N\}$, where $N$ denotes the number of video frames and $K$ is the number of keypoints. These trajectories are then visualized within video frames to enable further processing by the motion encoder, which shares the same architectural structure and initialization as the vision encoder. For camera motion understanding, inspired by [21], we introduce a camera encoder that processes a plücker embedding sequence $P\in\mathbb{R}^{N\times 6\times H\times W}$, where $H$, $W$ are the height and width of the video. This embedding accurately captures camera pose information, enabling precise modeling of camera trajectories. Finally, in line with Qwen2-VL, we employ special tokens to distinguish non-text conditions from texts. Besides the existing tokens, we introduce <|motion_start|>, <|motion_end|>, <|camera_start|>, <|camera_end|>, to demarcate the start and end of human and camera pose features.
@@ B-2 R-L METER BERTScore Accuracy Quality
To accurately interpret user generation intent under arbitrary conditions and yield structured target video captions, large-scale pretraining and instruction tuning are required. To this end, we adopt a two-stage training procedure: Stage-I: Alignment Learning. In this stage, as image and video encoders have been well-trained in Qwen2-VL, we only focus on aligning human pose features from the motion encoder and camera movement features with the word embeddings of the LLM backbone. To achieve this, we freeze the LLM and vision encoder while keeping the motion encoder trainable and optimizing it on a human pose description task. Similarly, for camera movement alignment, we unfreeze the camera encoder and train it on a camera movement description task, ensuring that camera-related conditions are embedded into the model’s latent space. This alignment phase establishes a strong foundation for effective representation learning for these conditions. Stage II: Condition-Interpreting Learning Building upon the aligned encoders and pretrained Qwen2-VL weights, we fine-tune the model on the Any2CapIns dataset for multimodal condition interpretation. However, direct fine-tuning leads to catastrophic forgetting due to the fixed output structure. To address this, we propose a progressive mixed training strategy. Specifically, the model is first trained on a single condition to establish a strong condition-specific understanding. As new conditions are introduced, we gradually incorporate vision-language instructions such as LLaVA-instruction [43] and Alpaca-52K [54]. This stepwise strategy ensures robust multimodal condition interpretation while preventing knowledge degradation.
@@ 5 Evaluation Suite
To assess caption generation quality from a lexical standpoint, we employ standard evaluation metrics commonly used in image/video captioning tasks, including BLEU [48], ROUGE [40], and METEOR [4]. We also introduce a Structural Integrity score to verify whether the generated captions adhere to the required six-component format, thereby ensuring completeness.
@@
To evaluate the semantic alignment of generated captions, we employ BERTScore [71], which computes similarity by summing the cosine similarities between token embeddings, effectively capturing both lexical and compositional meaning preservation. Additionally, we utilize CLIP Score [23] to assess the semantic consistency between the input visual condition and the generated videos.
@@
Evaluating structured captions typically focuses on their linguistic quality, yet it rarely assesses whether the model truly understands user intent and accurately reflects it across aspects such as style, emotion, and cinematic language. To bridge this gap and draw inspiration from [8], we introduce the Intent Reasoning Score (IRScore), a novel quantitative metric that leverages LLMs to assess whether generated captions accurately capture user intentions. IRScore identifies user-specified control factors, then decomposes the generated caption into QA pairs aligned with these factors. The framework has four steps: (1) User Intention Extraction: Analyze provided conditions to categorize user intent into six aspects: subject, background, movement, camera, interaction, and style. (2) Ground-Truth QA Pair Construction: Formulate aspect-specific QA pairs with defined requirements (e.g., for subject-related attributes, emphasize object count, appearance). (3) Predicted Answer Generation: Prompt GPT-4V to parse the predicted caption and generate answers based solely on it. (4) Answer Evaluation: Following [36], GPT-4V outputs two scores (correctness and quality) for each answer, which are then averaged across all QA pairs. More details are in Appendix §$\S$§E.
@@
The primary objective of generating structured captions is to enhance the video generation quality and controllability. Therefore, we adopt a series of metrics to assess the quality of videos generated based on structured captions. Following [26, 30], we evaluate video generation quality across four key dimensions: motion smoothness, dynamic degree, aesthetic quality, and image integrity. To further verify adherence to specific non-text conditions, we use specialized metrics: RotErr, TransErr, and CamMC [21] for camera motion accuracy; Mean Absolute Error (MAE) for depth consistency [18]; DINO-I [51], CLIP-I [51] Score to evaluate identity preservation under multiple identities, and Pose Accuracy (Pose Acc.) [45] to access the alignment in the generated videos.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Video serves as a fundamental medium for capturing real-world dynamics, making diverse and controllable video generation a key capability for modern artificial intelligence (AI) systems. Recently, video generation has gained significant attention, driven by advancements in Diffusion Transformers (DiT) [Open-Sora, DiVE-abs-2409-01595, Make-A-Video, Latte-abs-2401-03048, kuaishou], which have demonstrated the ability to generate realistic, long-duration videos from text prompts. These advancements have even led to industrial applications, such as filmmaking. However, we observe that a major bottleneck in the current video generation community lies in accurately interpreting user intention, so as to produce high-quality, controllable videos.
@@
In text-to-video (T2V) generation, studies [CogVideoX, MiraData-Ju00YWZX0S24, CogVideo-Hong0ZL023] have suggested that detailed prompts, specifying objects, actions, attributes, poses, camera movements, and style, significantly enhance both controllability and video quality. Thus, a series of works have explored video recaption techniques (e.g., ShareGPT4Video [ShareGPT4Video], MiraData [MiraData-Ju00YWZX0S24], and InstanceCap [InstanceCap]) to build dense structured captions for optimizing generative models. While dense captions are used during training, in real-world inference scenarios, users most likely provide concise or straightforward input prompts [InstanceCap]. Such a gap inevitably weakens instruction following and leads to suboptimal generation due to an incomplete understanding of user intent. To combat this, there are two possible solutions, manual prompt refinement or automatic prompt enrichment [InstanceCap, CogVideoX] using large language models (LLMs). Yet, these approaches either require substantial human effort or risk introducing noise from incorrect prompt interpretations. As a result, this limitation in precisely interpreting user intent hinders the adoption of controllable video generation for demanding applications such as anime creation and filmmaking.
@@
In addition, to achieve more fine-grained controllable video generation, one effective strategy is to provide additional visual conditions besides text prompts—such as reference images [Tune-A-Video-WuGWLGSHSQS23, I2V-Adapter-GuoZHGDWZLHZHM24], identity [abs-2411-17440, abs-2404-15275, abs-2402-09368], style [StyleMaster, StyleCrafter-LiuXZCXWWSY24], human pose [MaHCWC0C24, DreamPose-KarrasHWK23], or camera [CameraCtrl, CamI2V]—or even combinations of multiple conditions together [ControlVideo, Ctrl-Adapter, VideoComposer-WangYZCWZSZZ23]. This multimodal conditioning approach aligns well with real-world scenarios, as users quite prefer interactive ways to articulate their creative intent. Several studies have examined video generation under various conditions, such as VideoComposer [VideoComposer-WangYZCWZSZZ23], Ctrl-Adapter [Ctrl-Adapter], and ControlVideo [ControlVideo]. Unfortunately, these methods tend to rely on the internal encoders of Diffusion/DiT to parse rich heterogeneous input conditions with intricate requirements (e.g., multiple object IDs, and complex camera movements). Before generation, the model must accurately interpret the semantics of varied visual conditions in tandem with textual prompts. Yet even state-of-the-art (SoTA) DiT backbones have limited capacity for reasoning across different input modalities, resulting in suboptimal generation quality.
@@
To facilitate the any-to-caption instruction tuning for Any2Caption, we construct Any2CapIns, a large-scale dataset that converts a concise user prompt and diverse non-text conditions into detailed, structured captions. Concretely, the dataset encompasses four main categories of conditions: depth maps, multiple identities, human poses, and camera poses. Through extensive manual labeling combined with automated annotation by GPT-4V [Gpt-4v], followed by rigorous human verification, we curate a total of 337K high-quality instances, with 407K condition annotations, with the short prompts and structured captions averaging 55 and 231 words, respectively. In addition, we devise a comprehensive evaluation strategy to thoroughly measure the model’s capacity for interpreting user intent under these various conditions.
@@ 2 Related Work
Controllable video generation [abs-2411-04928, abs-2408-17424, abs-2411-18281, CameraCtrl] has long been a central topic in AI. Recent advanced DiT methods, such as OpenAI’s Sora [sora] and HunyuanVideo [kong2024hunyuanvideo], yield photorealistic videos over extended durations. Early work focused on text-controlled video generation [Make-A-Video, CogVideo-Hong0ZL023], the prevalent approach. Yet, text prompts alone may insufficiently capture user intent, spurring exploration of additional inputs including static images [Tune-A-Video-WuGWLGSHSQS23, I2V-Adapter-GuoZHGDWZLHZHM24], sketches [ControlVideo, VideoComposer-WangYZCWZSZZ23], human poses [PoseCrafter-ZhongZYYZL24, MaHCWC0C24, DreamPose-KarrasHWK23], camera views [CamI2V, CameraCtrl], and even extra videos [KaraKYRY24, DengWZTT24, I2VGen-XL]. Thus, unifying these diverse conditions into an “any-condition” framework is highly valuable.
Recent works such as VideoComposer [VideoComposer-WangYZCWZSZZ23], Ctrl-Adapter [Ctrl-Adapter], and ControlVideo [ControlVideo] have explored any-condition video generation. However, they face challenges in controlling multiple modalities due to the limited interpretability of text encoders in Diffusion or DiT. Motivated by existing MLLMs’ multimodal reasoning [llava-LiuLLL24, Video-LLaVA-LinYZCNJ024, Qwen2-VL], we propose leveraging an MLLM to consolidate all possible conditions into structured dense captions for better controllable generation. SoTA DiT models already exhibit the capacity to interpret dense textual descriptions as long as the input captions are sufficiently detailed in depicting both the scene and the intended generation goals. Thus, our MLLM-based encoder alleviates the comprehension bottleneck, enabling higher-quality video generation. To our knowledge, this is the first attempt in the field of any-condition video generation. Moreover, as the captioning stage is decoupled from backbone DiT, Any2Caption can integrate with existing video generation solutions without additional retraining.
Our approach also relates to video recaptioning, as our system produces dense captions from given conditions. In text-to-video settings, prior work [InstanceCap, CogVideoX, abs-2407-02371, IslamHYNTB24] shows that recaptioning yields detailed annotations that improve DiT training. For instance, ShareGPT4Video [ShareGPT4Video] uses GPT-4V [Gpt-4v] to reinterpret video content, while MiraData [MiraData-Ju00YWZX0S24] and InstanceCap [InstanceCap] focus on structured and instance-consistent recaptioning. Unlike these methods, we avoid retraining powerful DiT models with dense captions by training an MLLM as an any-condition encoder on pairs of short, dense captions that are easier to obtain. Moreover, recaptioning entire videos can introduce noise or hallucinations that undermine DiT training, whereas our framework sidesteps this risk. Finally, while previous studies rely on dense-caption-trained DiT models, the real-world user concise prompts might create a mismatch that degrades generation quality.
@@ 3 Any2CapIns Dataset Construction
While relevant studies recaption target videos for dense captions for enhanced T2V generation [ShareGPT4Video, InstanceCap, MiraData-Ju00YWZX0S24], these datasets exhibit two key limitations: (1) the absence of non-text conditions, and (2) short prompts that do not account for interactions among non-text conditions, potentially leading to discrepancies in real-world applications. To address these limitations, we introduce a new dataset, Any2CapIns, specifically designed to incorporate diverse multimodal conditions for generating structured video captions. The dataset is constructed through a three-step process (cf. Fig. 2), including data collection, structured caption construction, and user-centric short prompt generation.
@@
We begin by systematically categorizing conditions into four primary types: 1) Spatial-wise conditions, which focus on the structural and spatial properties of the video, e.g., depth maps, sketches, and video frames. 2) Action-wise conditions, which emphasize motion and human dynamics in the target video, e.g., human pose, motion. 3) Composition-wise conditions, which focus on scene composition, particularly in terms of object interactions and multiple identities in the target video. 4) Camera-wise conditions, which control video generation from a cinematographic perspective, e.g., camera angles, movement trajectories. Since it is infeasible to encompass all possible conditions in dataset collection, we curate representative datasets under each category, specifically including depth maps, human pose, multiple identities, and camera motion. During the data collection process, we leverage SoTA tools to construct conditions. For instance, Depth Anything [Depth-Anything-YangKH0XFZ24] is used to generate depth maps, DW-Pose [DW-Pose-YangZY023] provides human pose annotations, and SAM2 [SAM-2] is utilized for segmentation construction. In total, we collect 337K video instances and 407K conditions, with detailed statistics of the dataset presented in Tab. 1.
@@
The granularity of a caption, specifically the detailed selection of elements they encompass, plays a critical role in guiding the model to produce videos that closely align with the users’ intentions while preserving coherence and realism. Drawing inspiration from [MiraData-Ju00YWZX0S24], we inherit its structured caption format consisting of (1) Dense caption, (2) Main object caption, (3) Background caption, (4) Camera caption, and (5) Style caption. Furthermore, the action descriptions of the subjects significantly influence the motion smoothness of the videos [Koala-36M], we explicitly incorporate the (6) Action caption of subjects to form the final structured caption. Technically, following [Koala-36M], we independently generate each caption component and subsequently integrate them to construct the final structured caption.
@@ 4 Any2Caption Model
Similar to existing MLLMs [llava-LiuLLL24, NExT-GPT-Wu0Q0C24, Video-LLaVA-LinYZCNJ024], Any2Caption incorporates an image encoder $F_{\mathcal{I}}$, a video encoder $F_{\mathcal{V}}$, a motion encoder $F_{\mathcal{M}}$ and a camera encoder $F_{\mathcal{C}}\}$ to process non-text conditions. These encoders are then integrated into an LLM backbone $F_{\mathcal{LLM}}$ (i.e., Qwen2-LLM) to facilitate structured video captioning. Specifically, we leverage a ViT-based visual encoder from Qwen2-VL as $F_{\mathcal{I}}$ and $F_{\mathcal{V}}$ for the unified modeling of images and videos, achieving effective interpretation of input conditions represented in image or video formats, such as depth maps and multiple identities. To enable human pose understanding, we represent the extracted human pose trajectories as $\bm{H}$$\{(x_{n}^{k},y_{n}^{k})|k=1,\cdots,K,n=1,\cdots,N\}$, where $N$ denotes the number of video frames and $K$ is the number of keypoints. These trajectories are then visualized within video frames to enable further processing by the motion encoder, which shares the same architectural structure and initialization as the vision encoder. For camera motion understanding, inspired by [CameraCtrl], we introduce a camera encoder that processes a plücker embedding sequence $P\in\mathbb{R}^{N\times 6\times H\times W}$, where $H$, $W$ are the height and width of the video. This embedding accurately captures camera pose information, enabling precise modeling of camera trajectories. Finally, in line with Qwen2-VL, we employ special tokens to distinguish non-text conditions from texts. Besides the existing tokens, we introduce <|motion_start|>, <|motion_end|>, <|camera_start|>, <|camera_end|>, to demarcate the start and end of human and camera pose features.
@@ B-2 R-L METER BERTScore Accuracy Quality
To accurately interpret user generation intent under arbitrary conditions and yield structured target video captions, large-scale pretraining and instruction tuning are required. To this end, we adopt a two-stage training procedure: Stage-I: Alignment Learning. In this stage, as image and video encoders have been well-trained in Qwen2-VL, we only focus on aligning human pose features from the motion encoder and camera movement features with the word embeddings of the LLM backbone. To achieve this, we freeze the LLM and vision encoder while keeping the motion encoder trainable and optimizing it on a human pose description task. Similarly, for camera movement alignment, we unfreeze the camera encoder and train it on a camera movement description task, ensuring that camera-related conditions are embedded into the model’s latent space. This alignment phase establishes a strong foundation for effective representation learning for these conditions. Stage II: Condition-Interpreting Learning Building upon the aligned encoders and pretrained Qwen2-VL weights, we fine-tune the model on the Any2CapIns dataset for multimodal condition interpretation. However, direct fine-tuning leads to catastrophic forgetting due to the fixed output structure. To address this, we propose a progressive mixed training strategy. Specifically, the model is first trained on a single condition to establish a strong condition-specific understanding. As new conditions are introduced, we gradually incorporate vision-language instructions such as LLaVA-instruction [llava-LiuLLL24] and Alpaca-52K [alpaca]. This stepwise strategy ensures robust multimodal condition interpretation while preventing knowledge degradation.
@@ 5 Evaluation Suite
To assess caption generation quality from a lexical standpoint, we employ standard evaluation metrics commonly used in image/video captioning tasks, including BLEU [Papineni02bleu], ROUGE [lin-2004-rouge], and METEOR [banarjee2005]. We also introduce a Structural Integrity score to verify whether the generated captions adhere to the required six-component format, thereby ensuring completeness.
@@
To evaluate the semantic alignment of generated captions, we employ BERTScore [BERTScore], which computes similarity by summing the cosine similarities between token embeddings, effectively capturing both lexical and compositional meaning preservation. Additionally, we utilize CLIP Score [CLIPScore-HesselHFBC21] to assess the semantic consistency between the input visual condition and the generated videos.
@@
[…]
@@
The primary objective of generating structured captions is to enhance the video generation quality and controllability. Therefore, we adopt a series of metrics to assess the quality of videos generated based on structured captions. Following [VBench-HuangHYZS0Z0JCW24, MiraData-Ju00YWZX0S24], we evaluate video generation quality across four key dimensions: motion smoothness, dynamic degree, aesthetic quality, and image integrity. To further verify adherence to specific non-text conditions, we use specialized metrics: RotErr, TransErr, and CamMC [CameraCtrl] for camera motion accuracy; Mean Absolute Error (MAE) for depth consistency [SparseCtrl-GuoYRALD24]; DINO-I [DreamBooth-RuizLJPRA23], CLIP-I [DreamBooth-RuizLJPRA23] Score to evaluate identity preservation under multiple identities, and Pose Accuracy (Pose Acc.) [MaHCWC0C24] to access the alignment in the generated videos.